how much temperature does it take to boil 17 grams of ammonia in 500 grams of water


1
Expert's answer
2020-07-27T05:35:28-0400

Moles of NH3 = 17 / 17 = 1 mole .


Molality = moles / mass of solvent in kg .


Molality = 1 mol . / 0.5 kg .= 2 mol /kg . = m .


elevation in boilling point = (Kb ) * m ;


elevation in boilling point = ( 0.512 ) * 2 = 1.024 k .


boilling point of solution = ( 373 + 1.024 ) k .


So , boilling point of solution = 101.024 k . Answer .